# Understanding the Challenges and Promises of Developing Generative AI Apps: An Empirical Study

### v3

@@ 3.3. Multi-Stage User Review Refinement
    Prompt validation. To assess the generalizability of the final filtering prompt and reduce the risk of overfitting to specific review instances, we manually evaluate its performance on a separate set of previously unseen reviews of the same categories and the same size. The model demonstrates strong generalization capabilities, achieving an average classification accuracy of 90%.
@@
    Large-scale review filtering. We apply the validated filtering prompt to the full dataset to filter all non-informative reviews. Reviews are processed in batches of 100 to minimize omissions, as larger batches sometimes result in the model returning only a subset of reviews. Any unreturned reviews are consolidated into new batches and resent. This process reduces 920,824 reviews to 676,066 informative reviews, as shown in Table 2.
Our data collection, spanning September and October 2024, resulted in an initial dataset of 2,778,643 user reviews across 186 unique Gen-AI apps. Following our multi-stage cleaning process, we obtain a dataset of 676,066 informative reviews across 173 Gen-AI apps, which forms the basis for our subsequent analysis.
@@ 3.4. LLM-powered user review analysis
We use the refined user review data to conduct topic extraction, topic assignment, trend analysis, and manual analysis. For RQ1, we design and evaluate our topic extraction and assignment prompts. We experiment with different data noise levels (i.e., different stages) and numbers of few-shot examples. For RQ2, we extract the top 10 topics per app category and assign them to user reviews. We conduct manual thematic analysis for each Gen-AI topic to derive user insights from reviews. For RQ3, we analyze how Gen-AI topics evolve over time using clustering and manual review. In the following sections, we describe the experiment design and the findings for each RQ.
@@ 4. RQ1: How accurately can LLMs identify topics in user reviews of Gen-AI apps?
Effective topic extraction and assignment of user reviews are essential for understanding user needs and improving app features for Gen-AI apps. Prior work has shown that prompt engineering techniques can help tailor LLMs to specific tasks (Marvin_Hellen_Jjingo_Nakatumba-Nabende_2024; 10.1145/3736408). Applying LLMs to topic extraction and assignment requires well-designed prompts to achieve high accuracy. In our setup, we adopt this principle by refining prompts through multiple pilot runs on statistically significant samples to improve label consistency and clarity. In this RQ, we investigate two key aspects of prompt design: 1) the extent to which cleaning the input data prior to prompting improves accuracy, and 2) the effect of varying the number of few-shot examples in the prompt. The goal is to identify the optimal combination of preprocessing and few-shot-based prompting that enables more accurate and reliable topic extraction and assignment from user reviews.
@@ 4.1. Experiment Setup
To design an effective topic extraction prompt, we experiment with varying levels of data cleaning and different numbers of few-shot examples. As described in section 3, we design a four-stage data cleaning framework. Stage 0 includes basic textual cleaning. Stage 1 excludes reviews before the launch of ChatGPT in 2022 (Akhtar_2024). Stage 2 excludes short reviews with three words or less. Stage 3 excludes non-informative reviews using an LLM-based filtering approach. We conduct experiments on Stage 1, Stage 2, and Stage 3, but not on Stage 0. We exclude Stage 0 because while it performs essential preprocessing, it does not remove irrelevant data as done in Stage 1. We approach this problem in four steps: 1) we generate random samples of user reviews for topic extraction and assignment drawing from each stage of data cleaning, 2) we apply three prompting strategies with varying numbers of few-shot examples (i.e., 0-shot, 3-shot, and 5-shot) to extract topics from these samples, 3) assign the extracted topics to a new set of sample reviews using the same prompts, and (4) we evaluate the accuracy of the topic assignment to determine the optimal combination of data cleaning and few-shot examples. We summarize this process in Figure 4 and detail it in this section.
@@
Figure 5. Template of the topic extraction prompts for (E1: 0-shot, E2: 3-shot, and E3: 5-shot) used to extract the top topics from a large sample of stage $x$ and app category $y$. Insert category name is where we insert the name of the app category $y$. Insert sample of reviews is where we list our large sample of reviews separated by a new line escape.
Topic extraction. Because prior work reported mixed effects of few-shot examples (Wang_Yang_Wei_2023; Prakash_Wang_Hoang_Hee_Lee_2023), we design three experiments, i.e., E1: 0 shot, E2: 3 shot and E3: 5 shot, with three prompt configurations differing in the number of few-shot examples, i.e., 0-shot, 3-shot and 5-shot. These values are chosen based on prior research and practical considerations in prompt-based learning  (Dos_Santos_Oliveira_De_Jesus_Aljedaani_Eler_2023; Maram_LLM_Cure; Pham_Hoyle_Sun_Resnik_Iyyer_2024). In E1: 0 shot, the prompt contains only the instructions. In E2: 3 shot, the prompt includes the instructions with three shots, i.e., three examples of user reviews with their topics. In E3: 5 shot, the prompt includes the instructions with five examples of user reviews with their topics. For each cleaning stage $x\in\{\text{Stage 1},\text{Stage 2},\text{Stage 3}\}$ and app category $y$, we apply each of the three prompt configurations to the large sample $S_{\text{large}}(x,y)$. This results in a set of extracted top topics, denoted as $T(x,y,n)$, where $n\in\{0,3,5\}$ indicates the number of few-shot examples used in the prompt. Figure 5 shows the three prompt configuration templates.
To mitigate hallucination during topic extraction, we anchor the prompt in actual user reviews, instruct the model to output only ten distinct high-level topics that describe app features, functionalities, and utility, and provide few-shot examples that illustrate the expected level of abstraction and discourage free-form or overly creative topic invention. To further adapt the LLM to the Gen-AI domain, the few-shot examples are drawn from real reviews of Gen-AI apps, exposing the model to characteristic terminology and contexts (e.g., discussions of AI Performance, Technical Difficulties, and Content Quality).
Topic assignment. To evaluate the effectiveness of the extracted topic sets $T(x,y,n)$ produced by each topic extraction experiment with $n\in\{0,3,5\}$, we perform a follow-up topic assignment task using LLM. Specifically, for each experiment $n$, we design a dedicated assignment prompt, illustrated in Figure 6, that takes as input: 1) the category-specific set of top topics $T(x,y,n)$ extracted for that app category in that experiment, 2) five few-shot example reviews manually selected and labeled from the corresponding large sample $S_{\text{large}}(x,y)$, and 3) the target reviews from the small sample $S_{\text{small}}(x,y)$ to which topics are assigned. We manually select and assign topics to five few-shot examples to improve assignment accuracy as done in prior work (Maram_LLM_Cure; Deshmukh_Raut_Bhavsar_Gurav_Patil_2025).
@@
Accuracy evaluation. To assess the accuracy of the LLM-assigned topics, we conduct a manual annotation of the assignments. The first and second authors manually and independently label the correctness of LLM assignment. An assignment is considered correct if the LLM assigned topic accurately reflects the content of the review. The Cohen’s Kappa agreement score (cohen1960kappa) is computed, yielding a score of 0.56, indicative of moderate agreement. The two authors discuss and resolve the initial disagreements to align on labeling criteria. In total, 4,353 LLM topic assignments are manually evaluated for correctness for five categories, three stages, and three experiments. Table 3 shows the accuracy for each cleaning stage, app category and shot examples. To determine whether the observed accuracy differences are statistically significant, we perform statistical tests. Specifically, we apply the non-parametric test, Kruskal–Wallis H test (Chan_Walmsley_1997), to evaluate the effect of few-shot prompt configurations on overall accuracy. Additionally, post-hoc pairwise comparisons are conducted using the Mann–Whitney U test (Mann_Whitney_1947) with Bonferroni correction to adjust for multiple testing (3b6386c9-eeae-33a7-8c59-c0604aa87bef). We also use one-way ANOVA (Judd_McClelland_Ryan_2017) to test for significant differences across cleaning stages and few-shot settings.
@@
4.2. RQ1 Findings
Observation 1.1: Providing more examples significantly improves classification and assignment accuracy. The average accuracy increases across experiments for each cleaning stage, with E3: 5-shot consistently achieving the highest accuracy in all three stages, as shown in Table 3. Notably, E3: 5-shot is the only setting to achieve average accuracies above 80% across all samples. The Kruskal–Wallis test indicates that the number of few-shot examples has a statistically significant effect on overall accuracy ($H=7.778$, $p=0.0205$). Post hoc Mann–Whitney U tests with Bonferroni correction show that the accuracy improvement from E1: 0 shot to E3: 5 shot is statistically significant ($p=0.0181$), whereas differences between E1: 0 shot and E2: 3 shot ($p=1.000$) and E2: 3 shot and E3: 5 shot ($p=0.2519$) are not. These results suggest that increasing the number of few-shot examples, specifically 5-shot, significantly enhances overall model accuracy.
Observation 1.2: The proportion of informative reviews varies by cleaning stage, which impacts overall accuracy. As shown in Table 3, accuracy does not follow a consistent trend across the three cleaning stages (i.e., Stage 1, Stage 2, and Stage 3). For instance, the average accuracy for E1: 0 shot and E2: 3 shot increase from Stage 1 to Stage 2, but decrease from Stage 2 to Stage 3. In contrast, E3: 5 shot shows a slight decrease from Stage 1 to Stage 2, followed by a slight increase in Stage 3. This inconsistency may be explained by differences in the proportion of informative reviews across stages: 43% in Stage 1, 66% in Stage 2, and 88% in Stage 3. Since assigning topics to non-informative reviews is generally easier and less dependent on nuanced understanding, variations in their prevalence can skew overall accuracy. As a result, overall accuracy may not accurately reflect the model’s effectiveness in handling more meaningful and content-rich reviews across different cleaning stages.
@@ 4.2. RQ1 Findings
Observation 1.3: Stage 3 consistently yields the highest accuracy on informative reviews across all app categories and experiments. Given the inconsistencies observed in overall accuracy explained in observation 1.2, we isolate informative reviews to better evaluate model performance. Table 4 shows that, across all experiments and app categories, average accuracy on informative reviews is highest in Stage 3. This suggests that data cleaning improves the model’s ability to assign topics to reviews with informative content correctly. We conduct a one-way ANOVA test to evaluate whether the cleaning stage has a significant effect on classification accuracy. The ANOVA for few-shot experiments does not reveal significant differences ($F(2,42)=2.25$, $p=0.118$).
@@
Observation 1.4: The best performing setup combines filtering non-informative reviews, i.e., Stage 3 and more few shot examples, i.e., E3: 5-shot. Our findings suggest that using LLMs to filter the non-informative reviews and increasing the number of few-shot examples to five improves the accuracy of the model. Based on these findings, we adopt the Stage 3 – E3: 5-shot configuration for topic classification and assignment in RQ2, as it yielded the highest overall accuracy of 91% and the highest informative accuracy of 94%.
@@
Observation 1.5: Under the best-performing configuration (E3: 5-shot), only 9% of reviews were misclassified. Among these, 2% of all reviews involved incorrect topic assignments, 3% were missed topics (valid reviews marked as “Other”), and 4% were over-assignments (non-topical reviews given a label). Most errors occurred in short or ambiguous reviews with limited context, where topic boundaries were less clear. These findings indicate that misclassifications are infrequent and stem from understandable edge cases rather than uncontrolled generation or hallucination, reinforcing the reliability of the LLM-based labeling process.
@@
5. RQ2: What are the most prominent topics discussed in user reviews and in developer responses of Gen-AI apps?
Having validated the accuracy of our LLM-based topic extraction and assignment framework in RQ1, we now apply it to the full dataset to examine what users discuss in their reviews of Gen-AI apps. Understanding the issues that users care about is essential for improving the design and evaluation of Gen-AI apps. In RQ2, we uncover the top 10 topics that users discuss in reviews of Gen-AI apps across different app categories. We then analyze topic frequency and associated review ratings to interpret common user behaviours and needs, supporting the identification of actionable insights that inform the design of more adaptive and user-centered Gen-AI experiences.

### v4

@@ 3.3. Multi-Stage User Review Refinement
    Prompt validation. To assess the generalizability of the final filtering prompt and reduce the risk of overfitting to specific review instances, we manually evaluate its performance on a separate set of previously unseen reviews of the same categories and the same size. The model demonstrates strong generalization capabilities, achieving an average filtering accuracy of 90%.
@@
    Large-scale review filtering. We apply the validated filtering prompt to the full dataset to filter all non-informative reviews. Reviews are processed in batches of 100 to minimize omissions, as larger batches sometimes result in the model returning only a subset of reviews. Any unreturned reviews are consolidated into new batches and resent.
At the end of Stage 3, the dataset consists of informative reviews across 15 Google Play Store app categories. We then assess topic quality across categories and exclude eight categories with fewer than 1,000 reviews. Small sample sizes do not support the extraction of distinct topics, as reviews tend to overlap in content. Prior work shows that a fixed number of topics may be too high for some datasets, leading to over-clustering and insufficiently distinct topics (Greene et al., 2014).
To validate this threshold, we perform topic extraction across all app categories and observe that categories with fewer than 1,000 reviews consistently produce redundant and poorly differentiated topics. For example, in the Books & Reference category (342 informative reviews), multiple topics capture overlapping concepts, such as “Content Generation Features” and “Quality of Generated Content,” or “User Experience” and “User Interface.” A similar pattern is observed in the Health & Fitness category (729 reviews), where topics remain insufficiently distinct.
These redundancies introduce ambiguity during topic assignment, making it difficult to reliably determine which topic a review belongs to. Therefore, we adopt the 1,000-review threshold to ensure topic diversity, stability, and reliable labeling.
Our data collection, spanning October 2024 to February 2026, results in an initial dataset of 4,183,347 Google Play Store user reviews across 172 unique Gen-AI apps and 536,779 App Store reviews from 70 apps. Following our multi-stage cleaning process, we obtain a dataset of 1,035,342 informative Google Play Store reviews across 171 Gen-AI apps, and 225,469 App Store reviews from 59 apps, which form the basis for our subsequent analysis.
@@ 3.4. LLM-powered user review analysis
We use the refined user review data to conduct topic extraction, topic assignment, trend analysis, and manual analysis. For RQ1, we design and evaluate our topic extraction and assignment prompts. We experiment with different levels of data noise (i.e., different stages) and numbers of few-shot examples. For RQ2, we extract the top 10 topics per app category and assign them to user reviews. We conduct manual thematic analysis for each Gen-AI topic to identify challenges and opportunities derived from user review insights. For RQ3, we analyze how Gen-AI topics evolve over time using temporal clustering and manual review. In the following sections, we describe the experiment design and the findings for each RQ.
3.5. Adaptability of SARA Framework
In this study, SARA is empirically evaluated only on Gen-AI mobile apps. However, SARA is designed as a modular analysis framework in which the overall analytical framework remains fixed, while domain-specific configurations can be adapted. In particular, SARA can be applied to analyze user reviews of multiple mobile apps across different app categories and app stores by customizing domain-specific components. Because the framework operates on user-generated review data rather than Gen-AI-specific content, it can be adapted to analyze user perceptions of a wide range of mobile apps.
In the first stage, app selection is domain-agnostic: researchers may define and select target apps based on any domain-specific criteria or definitions relevant to their study. The second stage, review acquisition, follows the same procedure regardless of domain, as user reviews are collected from app store platforms using consistent metadata and retrieval mechanisms.
The review refinement stage is largely reusable across domains. Basic preprocessing steps (e.g., removal of white space, emojis, and short reviews) remain unchanged, and our findings in RQ1 show that LLMs are effective at filtering non-informative reviews using a general definition of informative reviews that is not tied to Gen-AI-specific functionality. As such, the non-informative review-filtering prompt can be applied without modification to other domains.
In the analysis stage, SARA employs topic extraction followed by topic assignment. While the overall prompt structure for topic extraction is reusable, the examples and topic labels must be adapted to reflect domain-specific concepts and functionalities. This customization is expected and enables the framework to capture contextually relevant themes. Similarly, topic assignment relies on a reusable prompt template, with domain-specific examples supplied by the researcher when applying the framework to a new domain.
Together, these design choices allow SARA to be adapted to multiple mobile apps across any app category and across different app stores by modifying domain-specific inputs and topic definitions, while preserving the framework’s core analytical stages and processing logic.
@@ 4. RQ1: How accurately can LLMs identify topics in user reviews of Gen-AI apps?
Effective topic extraction and assignment of user reviews are essential for understanding user needs and improving app features for Gen-AI apps. We identify the optimal combination of preprocessing and few-shot prompting to enable more accurate and reliable topic extraction and assignment from user reviews.
@@ 4.1. Experiment Setup
To design an effective topic extraction prompt, we conduct experiments on Stage 1, Stage 2, and Stage 3, but not on Stage 0. We exclude Stage 0 because, although it performs essential preprocessing, it does not remove irrelevant data as in Stage 1. We approach this problem in four steps: 1) we generate random samples of user reviews for topic extraction and assignment drawing from each stage of data cleaning, 2) we apply three prompting strategies with varying numbers of few-shot examples (i.e., 0-shot, 3-shot, and 5-shot) to extract topics from these samples, 3) assign the extracted topics to a new set of sample reviews using the same prompts, and (4) we evaluate the accuracy of the topic assignment to determine the optimal combination of data cleaning and few-shot examples. We summarize this process in Figure 3 and detail it in this section.
@@
Topic extraction. Because prior work reported mixed effects of few-shot examples (Wang et al., 2023; Prakash et al., 2023), we design three experiments, i.e., E1: 0 shot, E2: 3 shot and E3: 5 shot, with three prompt configurations differing in the number of few-shot examples, i.e., 0-shot, 3-shot and 5-shot. These values are chosen based on prior research and practical considerations in prompt-based learning  (Dos Santos et al., 2023; Assi et al., 2025a; Pham et al., 2024). In E1: 0 shot, the prompt contains only the instructions. In E2: 3 shot, the prompt includes the instructions with three shots, i.e., three examples of user reviews with their topics. In E3: 5 shot, the prompt includes the instructions with five examples of user reviews with their topics. For each cleaning stage $x\in\{\text{Stage 1},\text{Stage 2},\text{Stage 3}\}$ and app category $y$, we apply each of the three prompt configurations to the large sample $S_{\text{large}}(x,y)$. This results in a set of extracted top topics, denoted as $T(x,y,n)$, where $n\in\{0,3,5\}$ indicates the number of few-shot examples used in the prompt. The final version of the topic extraction prompt and the few-shot examples can be found in Appendix B.2.
To mitigate hallucinations during topic extraction, we anchor the prompt in actual user reviews, instruct the model to output only ten distinct high-level topics that describe app features, functionalities, and utility, and provide few-shot examples that illustrate the expected level of abstraction and discourage free-form or overly creative topic invention. To further adapt the LLM to the Gen-AI domain, the few-shot examples are drawn from real reviews of Gen-AI apps, exposing the model to characteristic terminology and contexts (e.g., discussions of AI Performance, Technical Difficulties, and Content Quality).
Topic assignment. To evaluate the effectiveness of the extracted topic sets $T(x,y,n)$ produced by each topic extraction experiment with $n\in\{0,3,5\}$, we perform a follow-up topic assignment task using LLM. Specifically, for each experiment $n$, we design a dedicated assignment prompt that takes as input: 1) the category-specific set of top topics $T(x,y,n)$ extracted for that app category in that experiment, 2) five few-shot example reviews manually selected and labeled from the corresponding large sample $S_{\text{large}}(x,y)$, and 3) the target reviews from the small sample $S_{\text{small}}(x,y)$ to which topics are assigned. We manually select and assign topics to five few-shot examples to improve assignment accuracy as done in prior work (Assi et al., 2025a; Deshmukh et al., 2025). The final topic assignment prompt is provided in Appendix B.3.
@@
Accuracy evaluation. To assess the accuracy of the LLM-assigned topics, we conduct a manual annotation of the assignments. The first and second authors manually and independently label the correctness of the LLM assignment. An assignment is considered correct if the LLM assigned topic accurately reflects the content of the review. The Cohen’s Kappa agreement score (Cohen, 1960) is computed, yielding a score of 0.56, indicative of moderate agreement. The two authors discuss and resolve the initial disagreements to align on labeling criteria.
We analyze cases of annotator disagreement to better understand ambiguity in topic classification. Disagreements primarily occur in reviews with limited context or broad, non-informative statements (e.g., the “brief expressions of satisfaction” topic identified in earlier stages), where the relevance of a specific topic is unclear. In these cases, annotators differ in their strictness: one may accept the assigned topic as valid, while another may judge the review as too vague or incoherent to support that assignment. Upon discussion, we label such assignments as incorrect, as the reviews do not provide sufficient or interpretable evidence for the assigned topic, and therefore should have been marked as “other” by the LLM. These disagreements highlight the challenges of topic classification in earlier stages, where noisy and non-informative reviews can lead to the extraction of generic topics. As a result, assigning topics becomes difficult due to overlapping themes and the absence of clear semantic signals in the underlying reviews.
In total, 4,353 LLM topic assignments are manually evaluated for correctness for five categories, three stages, and three experiments. Table 3 shows the accuracy for each cleaning stage, app category and shot examples. We supplement this with Table 4, which reports accuracy on informative reviews across cleaning stages, app categories, and shot settings. This separation enables a fair comparison, as non-informative reviews are easier to classify.
@@
To determine whether the observed accuracy differences are statistically significant, we perform statistical tests. Specifically, we apply the non-parametric test, Kruskal-Wallis H test (Chan and Walmsley, 1997), to evaluate the effect of few-shot prompt configurations on overall accuracy. Additionally, post-hoc pairwise comparisons are conducted using the Mann-Whitney U test (Mann and Whitney, 1947) with Bonferroni correction to adjust for multiple testing (Cabin and Mitchell, 2000). We also use one-way ANOVA (Judd et al., 2017) to test for significant differences across cleaning stages and few-shot settings.
4.2. RQ1 Findings
Finding 1.1: Providing more examples significantly improves classification and assignment accuracy. The average accuracy increases across experiments for each cleaning stage, with E3: 5-shot consistently achieving the highest accuracy in all three stages, as shown in Table 3. Notably, E3: 5-shot is the only setting to achieve average accuracies above 80% across all samples. The Kruskal-Wallis test indicates that the number of few-shot examples has a statistically significant effect on overall accuracy ($H=7.778$, $p=0.0205$). Post hoc Mann-Whitney U tests with Bonferroni correction show that the accuracy improvement from E1: 0 shot to E3: 5 shot is statistically significant ($p=0.0181$), whereas differences between E1: 0 shot and E2: 3 shot ($p=1.000$) and E2: 3 shot and E3: 5 shot ($p=0.2519$) are not. These results suggest that increasing the number of few-shot examples, specifically 5-shot, significantly enhances overall model accuracy.
Finding 1.2: The proportion of informative reviews varies by cleaning stage, which impacts overall accuracy. As shown in Table 3, accuracy does not follow a consistent trend across the three cleaning stages (i.e., Stage 1, Stage 2, and Stage 3). For instance, the average accuracy for E1: 0 shot and E2: 3 shot increase from Stage 1 to Stage 2, but decrease from Stage 2 to Stage 3. In contrast, E3: 5 shot shows a slight decrease from Stage 1 to Stage 2, followed by a slight increase in Stage 3. This inconsistency may be explained by differences in the proportion of informative reviews across stages: 43% in Stage 1, 66% in Stage 2, and 88% in Stage 3. Since assigning topics to non-informative reviews is generally easier and less dependent on nuanced understanding, variations in their prevalence can skew overall accuracy. As a result, overall accuracy may not accurately reflect the model’s effectiveness in handling more meaningful and content-rich reviews across different cleaning stages.
@@ 4.2. RQ1 Findings
Finding 1.3: Stage 3 consistently yields the highest accuracy on informative reviews across all app categories and experiments. Given the inconsistencies observed in overall accuracy explained in finding 1.2, we isolate informative reviews to better evaluate model performance. Table 4 shows that, across all experiments and app categories, average accuracy on informative reviews is highest in Stage 3. This suggests that data cleaning improves the model’s ability to assign topics to reviews with informative content correctly. We conduct a one-way ANOVA test to evaluate whether the cleaning stage has a significant effect on classification accuracy. The ANOVA for few-shot experiments does not reveal significant differences ($F(2,42)=2.25$, $p=0.118$).
@@
Finding 1.4: The best performing setup combines filtering non-informative reviews, i.e., Stage 3 and more few shot examples, i.e., E3: 5-shot. Our findings suggest that using LLMs to filter the non-informative reviews and increasing the number of few-shot examples to five improves the accuracy of the model. Based on these findings, we adopt the Stage 3 - E3: 5-shot configuration for topic classification and assignment in RQ2, as it yielded the highest overall accuracy of 91% and the highest informative accuracy of 94%.
@@
Finding 1.5: Under the best-performing configuration (E3: 5-shot), only 9% of reviews were misclassified. Among these, 2% of all reviews involved incorrect topic assignments, 3% were missed topics (valid reviews marked as “Other”), and 4% were over-assignments (non-topical reviews given a label). Most errors occurred in short or ambiguous reviews with limited context, where topic boundaries were less clear. These findings indicate that misclassifications are infrequent and stem from understandable edge cases rather than uncontrolled generation or hallucination, reinforcing the reliability of the LLM-based labeling process.
@@
5. RQ2: What are the most prominent topics discussed in user reviews?
We identify the top 10 topics discussed in user reviews across app categories and analyze their frequency and associated ratings. This analysis uncovers common user concerns, expectations, and usage patterns, from which we derive key opportunities, challenges, and actionable insights for designing more effective, user-centered Gen-AI systems.